OpenStudy (clara1223):

find the limit at x approaches -5 of (30-6x)/(x+5) a) 12 b) 6 c) does not exist d) 0 e) -6/5

OpenStudy (clara1223):

I already took a 6 out of the numerator so i have 6(-x+5)/(x+5)

OpenStudy (lynfran):

\[\lim_{x \rightarrow -5}\frac{ (30-6x) }{ (x+5) }\]\[\lim_{x \rightarrow -5}\frac{ 6(5-x) }{ (x+5) }\]\[\lim_{x \rightarrow -5}\frac{ 6(5-(-5)) }{ (-5+5) }\]

OpenStudy (lynfran):

u would get zero in the denominator so the limit DNE

OpenStudy (clara1223):

But then you have a 0 in the denominator. So does the limit not exist?

OpenStudy (clara1223):

Oh, thank you

OpenStudy (anonymous):

Why isnt infinity included in your choices?

OpenStudy (lynfran):

welcome

OpenStudy (clara1223):

@joyraheb another way to say infinity or negative infinity is that the limit doesnt exist